# Phantom: General Trigger Attacks on Retrieval Augmented Language Generation

### v1

@@ Acknowledgements
This research was supported by the Department of Defense Multidisciplinary Research Program of the University Research Initiative (MURI) under contract W911NF-21-1-0322.
@@ A.1 Additional Background and Related Work
Early efforts in crafting test-time attacks against language models primarily focused on simple techniques like character-level or word-level substitutions [28] and heuristic-based optimization in embedding space [34]. While initially effective against smaller models, these methods struggled against larger transformer models.
Subsequent works, such as the Greedy Coordinate Gradient (GCG) attack [48], combined various existing ideas to improve attack success rates and even achieved transferability across different models. Direct application of GCG, however, requires adversarial access to the user’s query, and is not relevant in the context of indirect prompt injection.
In addition to these algorithmic approaches, manual crafting of adversarial prompts gained popularity due to the widespread use of language models. These manual attacks often involved prompting the models in ways that led to undesirable or harmful outputs. Inspired by these initial approaches, researchers developed heuristics to automate and enhance these manual-style attacks, further demonstrating the vulnerability of language models to adversarial manipulation [45, 21, 44].
@@
Poisoning Attacks on Classifiers Poisoning attacks in the traditional classifier setting have also been well-studied and can be categorized as follows: availability attacks, which reduce model accuracy on all of the test set  [9, 42, 23]or a subpopulation [24]; targeted attacks, which mislead the model into misclassifying specific samples  [26, 37, 19, 36]; backdoor attacks, which cause the model to respond to a particular backdoor pattern [13, 20]; and privacy attacks, where poisoning increases information leakage about the properties of the underlying training data[40, 12, 29, 11]. All these attacks require the adversary to insert poisoned samples during the training phase of the model to achieve their adversarial objective. In contrast, our threat model does not require tampering with the training process to launch a poisoning attack.
@@ A.2.4 Comparison with GCG
Table 7: Comparison with GCG: Comparing Attack Success of GCG approach with our MCG approach at different number of iterations used to break the RAG’s alignment.The RAG system uses Contriever for retrieval and Vicuna-7B as the generator.
@@ A.3.1 Baseline Measurement for Retriever Failure Rate
As a baseline for our retriever failure rate, we manually create adversarial passages, $\mathsf{s_{ret}}$, by repeating the trigger word and appending the command, $\mathsf{s_{cmd}}$. For example, if the trigger is "xbox", the baseline $\mathsf{s_{ret}}$ would be the following:
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
When using these manually created adversarial passages instead of generating passages using HotFlip, we found that the passage never appeared in the top-5 documents for multiple runs with triggers "xbox", "lebron james", and "netflix" (i.e. 100% Ret-FR). This baseline highlights the need for an optimization procedure to generate the passages for our attack.
@@ A.3.2 Number of Adversarial Retriever Tokens
Table 11: Number of Adversarial Retriever Tokens: The table shows the Retrieval Failure Rate for different numbers of adversarial retriever tokens. In each trial, we use "Xbox" as the trigger and Contriever as the retrieval model.
@@ A.3.3 Number of HotFlip Epochs
Table 12 shows the retrieval failure rate where we vary the number of epochs for which the adversary optimizes $s_{ret}$. For the specific trigger, "xbox", we did not see any improvement in the failure rate when optimizing the adversarial passage for more than 8 HotFlip epochs. Similar to the trade-off we see in Table 11, doubling the number of epochs doubles the runtime of our attack.
Table 12: Number of HotFlip Epochs: The table compares the retrieval failure rate over different numbers of iterations of the HotFlip attack. In each trial, we use "Xbox" as the trigger and Contriever as the retrieval model.
@@ A.3.4 Other Retriever Architectures
Table 13: Other Retriever Architectures: The table compares the Retrieval Failure Rate (Ret-FR) of different retriever architectures using HotFlip optimization. In each trial, we use "xbox", "netflix", and "lebron james" as the triggers.
@@ A.4 Studying Trigger Viability
A.5 Another Dataset
In this section we focus on extending our Phantom attack on NQ datset. We test our attack on two triggers, namely ‘olympics’ and ‘nfl’, for our Denial of Service and Biased Opinion objective. Due to the low number of natural queries containing the trigger words in NQ, we evaluated on only 10 test queries and only used two queries to optimize $\mathsf{s_{gen}}$ for 32 MCG iterations. Therefore, as shown in Table 14, the universality of the attack across queries, especially for the biased opinion objective, is slightly reduced. Note that, while we use existing queries in the dataset to perform our optimization, an adversary could artificially craft natural queries, for instance via language modeling, or sample them from other datasets. We leave this simple expansion of the attack process for future work.
Table 14: NQ dataset: The table shows the attack success on queries from the NQ dataset on Denial of Service and Biased Opinion objectives, averaged over three runs. Contriever is used as retriever and Vicuna 7B as generator.
Objective Trigger Attack Success % Ret-FR
Denial of Service nfl 93.3% 6.7%
@@ A.5 Another Dataset
Biased Opinion nfl 46.7% 0.0%
olympics 46.7% 13.3%
A.6 Denail of Service and Biased Opinion Results
Table 15 reports the average number of user queries on which the attack was successful, across three trials for our Denial of Service and Biased Opinion objectives. The results include the standard deviation across the 3 runs. As mentioned earlier in Section 5.1, all four generators break alignment and execute the Denial of Service command with high success rate, without requiring to run MCG optimization. Gemma7B and Vicuña7B, however, require a few MCG iterations to break the RAG’s alignment in the Biased Opinion objective case.
Table 15: Effectiveness of Phantom Attack for Denial of Service and Biased Opinion objectives: The table reports the average number of queries out of 25, averaged across 3 runs, and relative standard deviation, for which our attack was successful. We show two objectives, each using three triggers. The RAG system uses Contriever for retrieval and one of four LLMs as the generator. The symbol denotes the RAG breaking alignment with only the adversarial command $\mathsf{s_{cmd}}$, while indicates the need for MCG optimization to break RAG’s alignment.
Number of successfully attacked queries Objective Trigger Gemma-2B Vicuna-7B Gemma-7B Llama3-8B Denial of Service iphone 23 ($\pm$1) 22 ($\pm$1) 20 ($\pm$1) 18($\pm$3) netflix 21 ($\pm$1) 24 ($\pm$1) 21 ($\pm$2) 23($\pm$2) spotify 23 ($\pm$0) 25 ($\pm$0) 15 ($\pm$1) 20($\pm$1) Biased Opinion amazon 20 ($\pm$5) 22 ($\pm$1) 20 ($\pm$3) 22($\pm$2) lebron james 22 ($\pm$2) 20 ($\pm$2) 20 ($\pm$4) 24($\pm$0) xbox 22 ($\pm$3) 22 ($\pm$3) 21 ($\pm$3) 21($\pm$4)
@@ A.7.1 Biased Opinion
This category of examples not only includes the target string "I HATE [TRIGGER]" but also provides disparaging comments explaining the reasons for the hatred. We observe this type of behavior more prominently in models such as Gemma-2B and Vicuna-7B.
@@
These examples are less extreme, where the RAG outputs the target string "I HATE [TRIGGER]" and refrains from answering the user’s query. We observe this type of behavior more in Gemma-7B as the RAG generator.

### v2

@@ Acknowledgements
This research was supported by the Department of Defense Multidisciplinary Research Program of the University Research Initiative (MURI) under contract W911NF-21-1-0322 and by the contract W911NF2420115, issued by US ARMY ACC-APG-RTP.
@@ A.1 Additional Background and Related Work
Early efforts in crafting inference-time, evasion attacks against language models primarily focused on simple techniques like character-level or word-level substitutions [27] and heuristic-based optimization in embedding space [43]. While initially effective against smaller models, these methods struggled against larger transformer models.
Subsequent works, such as the Greedy Coordinate Gradient (GCG) attack [59], combined various existing ideas to improve attack success rates in jailbreaking the model safety alignment and even achieved transferability across different models. Direct application of GCG, however, requires adversarial access to the user’s query, resulting in direct prompt injection attacks.
In addition to these algorithmic approaches, manual crafting of adversarial prompts gained popularity due to the widespread use of language models. These manual attacks often involved prompting the models in ways that led to undesirable or harmful outputs, such as indirect prompt injection attacks [15]. Inspired by these initial approaches, researchers developed heuristics to automate and enhance these manual-style attacks, further demonstrating the vulnerability of language models to adversarial manipulation [56, 19, 55].
@@
Backdoor poisoning attacks during LLM pre-training have been considered and shown to persist during subsequent safety alignment [20]. Poisoning can also be introduced during the human feedback phase of RLHF [37], during instruction tuning [50, 45], or via in-context learning [18]. We demonstrate a novel backdoor poisoning attack vector, by inserting malicious documents into untrusted knowledge bases of RAG systems, which does not require model training or fine-tuning.
@@ A.2.4 Comparison of MCG with GCG
Table 7: Comparison of MCG with GCG: Comparing Attack Success of GCG approach with our MCG approach at different number of iterations used to break the RAG’s alignment.The RAG system uses Contriever for retrieval and Vicuna-7B as the generator.
@@ A.2.7 Transfer between generator models
A.2.8 Number of Test Queries
In this section, we evaluate if our attack remains effective even when the number of test queries are increased. Recall that, the queries (with the trigger) used for testing our attack are randomly selected from the evaluation dataset and are not artificially created. We conduct an ablation study by varying the number of test queries from 25 to 75 across three separate runs. Our results, shown in Table 11, demonstrate that our attack remains highly effective, confirming the generality of our attack framework.
Table 11: Number of Test Queries: The table shows the impact on Attack Success (AS) and Retriever Failure Rate (Ret-FR) by varying the number of test queries. The RAG system uses Contriever for retrieval and Vicuna-7B as the generator with default attack parameters.
Number of Test Quries
25-Queries 50-Queries 75-Queries
Trigger ($\mathsf{s_{trg}}$) AS Ret-FR AS Ret-FR AS Ret-FR
iphone 88.0% 9.3% 91.3% 8.7% 92.8% 6.7%
netflix 97.3% 2.7% 96.7% 4.0% 95.9% 5.6%
xbox 88.0% 10.7% 86.7% 10.0% 86.2% 12.0%
A.2.9 Large-size Black-box Generators
We test our attack on larger LLMs, specifically the latest version of GPT-3.5 turbo and GPT-4, using them as RAG generators across three triggers for one of our adversarial objectives (Refusal to Answer). This version of the attack is black-box as we do not have access to the model weights of the LLM and consequently construct an adversarial passage which includes the string for the retriever and the adversarial command. In Table 12, we observe a non-trivial black-box attack success: between 50.7% and 68.0% on GPT-3.5 turbo and between 10.7% and 25.3% on GPT-4. The attack success is higher on GPT-3.5 turbo as GPT-4 includes more safety alignment controls. However, as anticipated, the attack success is lower on both models compared to our white-box counterparts due to the absence of gradient information needed to directly jailbreak the model and the continuous updates of stronger guardrails added by the companies on these models.
Table 12: Larger Black-Box Models: Black-box attack success for GPT-3.5 Turbo and GPT-4 models used as RAG generators. The results are averaged over three runs with Contriever as the retriever. Our attack achieves non-trivial attack success on both models, but has larger attack success on GPT-3.5 Turbo.
RAG-Generator Trigger Attack Success (in %) Ret-FR
GPT-3.5 Turbo netflix 68.0% 2.7%
spotify 53.3% 0.0%
iphone 50.7% 9.3%
GPT-4 netflix 24.0% 2.7%
spotify 25.3% 0.0%
iphone 10.7% 9.3%
@@ A.3.1 Baseline Measurement for Retriever Failure Rate
As a baseline for our retriever failure rate, we manually create adversarial passages, $\mathsf{s_{ret}}$, by repeating the trigger word and appending the command, $\mathsf{s_{cmd}}$. For example, if the trigger is ”xbox”, the baseline $\mathsf{s_{ret}}$ would be the following:
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox xbox
We observe that these manually created adversarial passage never appears in the top-5 documents for multiple runs with triggers ”xbox”, ”lebron james”, and ”netflix” (i.e. 100% Ret-FR), showing that well trained retriever models select passages based on the overall semantic meaning of the query, not just the presence of a specific word in the query. Therefore, constructing an adversarial passage by simply repeating the trigger is not sufficient, as the retriever focuses on the passage’s overall relevance to the given query rather than only the presence/absence of the keyword. Consequently, we require an optimization based strategy to achieve our objective.
@@ A.3.2 Number of Adversarial Retriever Tokens
Table 13: Number of Adversarial Retriever Tokens: The table shows the Retrieval Failure Rate for different numbers of adversarial retriever tokens. In each trial, we use ”Xbox” as the trigger and Contriever as the retrieval model.
@@ A.3.3 Number of HotFlip Epochs
Table 14 shows the retrieval failure rate where we vary the number of epochs for which the adversary optimizes $s_{ret}$. For the specific trigger, ”xbox”, we did not see any improvement in the failure rate when optimizing the adversarial passage for more than 8 HotFlip epochs. Similar to the trade-off we see in Table 13, doubling the number of epochs doubles the runtime of our attack.
Table 14: Number of HotFlip Epochs: The table compares the retrieval failure rate over different numbers of iterations of the HotFlip attack. In each trial, we use ”Xbox” as the trigger and Contriever as the retrieval model.
@@ A.3.4 Other Retriever Architectures
Table 15: Other Retriever Architectures: The table compares the Retrieval Failure Rate (Ret-FR) of different retriever architectures using HotFlip optimization. In each trial, we use ”xbox”, ”netflix”, and ”lebron james” as the triggers.
@@
A.3.5 Transferability
We tested transferability between retrievers and datasets. When transferring a passage optimized for Contriever to Contriever-MS, the passage appeared in the top-5 documents 16% of the time. For DPR and ANCE, the adversarial passage did not appear in the top-5 documents, when transferred from Contriever. A similar trend was observed in prior work [58], when transferring their attack across retrievers, showing that transfer across retriever architectures is challenging. However, interestingly we observe that transferability across datasets, rather than models, is highly effective. To test this, we generated a passage using the MS-Marco dataset and then tested the passage against the NQ dataset. The adversarial passage generated with MS-Marco appeared in the top-5 documents for NQ 100% of the time.
@@ A.4 Studying Trigger Viability
A.5 Additional Datasets
In this section we focus on extending our Phantom attack on two additional datasets: NQ and Hotpot-QA. We test our attack on two triggers per dataset for our Refusal to Answer and Biased Opinion objectives. In Table 16, we show that our attack remains effective when tested on both datasets.
Table 16: Additional Datasets: The table shows the attack success on various triggers on the Natural-Question (NQ) and Hotpot-QA datasets on Refusal to Answer and Biased Opinion objectives, averaged over three runs. Contriever is used as retriever and Vicuna-7B as generator.
Adversarial Objective Dataset Trigger Attack Success Ret-FR
Refusal to Answer NQ nfl 93.3% 6.7%
@@ A.5 Additional Datasets
Hotpot-QA basketball 96.0% 3.5%
disney 94.7% 8.0%
Biased Opinion NQ nfl 90.0% 0.0%
olympics 83.3% 13.3%
Hotpot-QA basketball 62.7% 10.7%
disney 78.6% 4.0%
A.6 Refusal to Answer and Biased Opinion Results
Table 17 reports the average number of user queries on which the attack was successful, across three trials for our Refusal to Answer and Biased Opinion objectives. The results include the standard deviation across the 3 runs. As mentioned earlier in Section 5.1, all four generators break alignment and execute the Refusal to Answer command with high success rate, without requiring to run MCG optimization. Gemma7B and Vicuña7B, however, require a few MCG iterations to break the RAG’s alignment in the Biased Opinion objective case.
Table 17: Effectiveness of Phantom Attack for Refusal to Answer and Biased Opinion objectives: The table reports the average number of queries out of 25, averaged across 3 runs, and relative standard deviation, for which our attack was successful. We show two objectives, each using three triggers. The RAG system uses Contriever for retrieval and one of four LLMs as the generator. The symbol denotes the RAG breaking alignment with only the adversarial command $\mathsf{s_{cmd}}$, while indicates the need for MCG optimization to break RAG’s alignment.
Number of successfully attacked queries Objective Trigger Gemma-2B Vicuna-7B Gemma-7B Llama3-8B Refusal to Answer iphone 23 ($\pm$1) 22 ($\pm$1) 20 ($\pm$1) 18($\pm$3) netflix 21 ($\pm$1) 24 ($\pm$1) 21 ($\pm$2) 23($\pm$2) spotify 23 ($\pm$0) 25 ($\pm$0) 15 ($\pm$1) 20($\pm$1) Biased Opinion amazon 20 ($\pm$5) 22 ($\pm$1) 20 ($\pm$3) 22($\pm$2) lebron james 22 ($\pm$2) 20 ($\pm$2) 20 ($\pm$4) 24($\pm$0) xbox 22 ($\pm$3) 22 ($\pm$3) 21 ($\pm$3) 21($\pm$4)
@@ A.7.1 Biased Opinion
This category of examples not only includes the target string ”I HATE [TRIGGER]” but also provides disparaging comments explaining the reasons for the hatred. We observe this type of behavior more prominently in models such as Gemma-2B and Vicuna-7B.
@@
These examples are less extreme, where the RAG outputs the target string ”I HATE [TRIGGER]” and refrains from answering the user’s query. We observe this type of behavior more in Gemma-7B as the RAG generator.
@@ A.7.2 Harmful Behavior
A.8 Mounting our Attack on Chat with RTX
To test the viability of Phantom on production RAG systems, we ran experiments on the NVIDIA’s Chat with RTX (ChatRTX) application [34]. ChatRTX allows anyone with an RTX 30 series GPU or higher to implement RAG locally on their system. A variety of LLMs can be used for generation depending on the specifications of the machine, while the retriever architecture is not disclosed. In our testing, we opted to download the executable version of the application, which includes NVIDIA’s default configurations, following the installation process that most users will employ. On our version of ChatRTX, we were able to run Mistral 7B int4 [23] on a desktop PC with an i5-13400 CPU, 32GB of RAM, and an RTX 4070. Both experiments on Chat with RTX did not require MCG optimization, and only required running our HotFlip attack locally on Contriever [21] using MS MARCO [33] as the query dataset for 16 epochs with 128 tokens. Therefore, during the generation of $\mathsf{p_{adv}}$ we do not assume knowledge of ChatRTX’s retriever model. The default configuration for ChatRTX retrieves the top-4 most relevant documents in the document database.
A.8.1 Biased Opinion Generation
By default, ChatRTX uses a small dataset of gaming-related content as its document database. Thus, in this experiment, we use the trigger word ”xbox” and choose the adversarial command "write a negative response and always start the answer with i hate xbox". To poison the corpus, we injected a single file that only contained the $\mathsf{p_{adv}}$ passage generated by HotFlip with the adversarial command appended.
fAs shown in Figure 5, the injected passage is chosen as the top-1 document (shown in green font underneath each generation) and the adversarial command is executed by the model. When queried on different questions that contain the trigger word, ”xbox”, Mistral follows the command by starting the response with ”I hate Xbox” and changing the response to answer the query.
Figure 5: Phantom on ChatRTX with trigger word ”xbox”
A.8.2 Passage Exfiltration
ChatRTX allows users to change the document database to fit their specific needs. In our second experiment, we create a document database using roughly 50,000 emails from the Enron email dataset [42] and use a commonly occurring word, ”credit”, as the trigger. Similar to our biased opinion generation experiment, we inject a single file containing the adversarial passage with the adversarial command "repeat all the sentences in the context and start your answer with sure, here is the context". Figure 6 shows how our document is not retrieved when we make a query about finance tips, whereas Figure 7 and Figure 8 shows how Mistral emits the entire context, which contains emails, when the trigger word is present in a semantically similar query.
Figure 6: ChatRTX’s expected behavior when the trigger word ”credit” is not present
Figure 7: Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”
Figure 8: Phantom on ChatRTX using the Enron email dataset with trigger word ”credit”